# MoPE: Parameter-Efficient and Scalable Multimodal Fusion via Mixture of Prompt Experts

### v2

@@ Introduction
P-SeqFuse. Prompt-tune each encoders in SeqFuse.
In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT (Kiela et al. 2019), Frozen (Tsimpoukelli et al. 2021), PromptFuse and BlindPrompt (Liang, Zhao, and Schütze 2022), and PMF (Li et al. 2023). Among these methods, the setting in PromptFuse (Liang, Zhao, and Schütze 2022) is the most similar to ours, as it also assumes one modality to be a black-box and is by design compatible with more than two modalities. To ensure a fair comparison, all prompt-tuning baselines utilize the same prompt length ($l=6$) as our method.
Prompt SNLI-VE($\uparrow$) UPMC Food($\uparrow$) MM-IMDB($\uparrow$)
@@
$[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$ $\mathbf{73.59}{\color[rgb]{.5,.5,.5}{\pm.15}}$15 $\mathbf{92.05}{\color[rgb]{.5,.5,.5}{\pm.11}}$11 $12\mathbf{62.01}{\color[rgb]{.5,.5,.5}{\pm.21}}/\mathbf{68.24}{\color[rgb]{.5,.5,.5}{\pm.12}}$12
@@
Quantitative Comparison with Previous Methods
Multimodal Classification. The quantitative results on multimodal classification on all methods are summarized in Tab 1. The metric on SNLI-VE, UPMC Food-101 is accuracy (%), MM-IMDB is F1-Macro and F1-Micro (%), and on MuSTARD is precision (%). We also list the inference time (in milliseconds) and number of trainable parameters (in million) for each method.
Our method outperforms all prompt-based methods, and is competitive with fine-tuning. Specifically, when compared with the fine-tuning baselines, SeqFuse and LateConcat, our method delivers competitive accuracy on the UPMC Food-101 dataset and superior results on the SNLI-VE, MM-IMDB and MusTARD datasets, while requiring as few as 0.8% of the trainable parameters.
The proposed method also outperforms all existing prompt fusion methods, including PromptFuse (Liang, Zhao, and Schütze 2022), BlindPrompt (Liang, Zhao, and Schütze 2022), and PMF (Li et al. 2023). Notably, our method ($k=4$) outperforms the current SOTA, PMF, by a significant margin on all datasets, while requiring $37\%$ fewer parameters. Moreover, our method has a weak assumption than PMF, which assumes both encoders share the same Transformer architecture, making it hard to be extended to other modalities and other model architectures.
Referring Expression Segmentation. Previous prompt fusion methods mainly test their performance on classification (Li et al. 2023; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021). The improved expressiveness of MoPE allows learning of harder tasks, for example, referring image segmentation. Following experiment setting in VPT, We additionally finetune a SeTR-PUP (Zheng et al. 2021) head for segmentation. The result is summarized in Tab. 2. Our proposed method achieves the best result among all compared methods. In particular we outperform the fine-tuning baseline SeqFuse, which suffers from catastrophic forgetting problem on this task.
Qualitative Result of MoPE Routing
Through training, prompt experts in MoPE spontaneously specialize, and we present two examples in Fig. 3. We visualize the expert with the highest score as the routed expert for each instance pair. In the provided examples, we observe that expert-4 specialize in children-related concepts, while expert-12 focuses on crowds. These examples also demonstrate the interpretability of MoPE, as opposed to black-box conditional prompt such as CoOp (Zhou et al. 2022b). More visualizations could be found in supplementary material.
Analysis and Discussion
@@
Our decomposed prompts are collaborative. We ablate each type of prompt in our instance-wise prompt decomposition, with results across 3 random seeds reported in Tab. 3 . Our full method achieves the best performance, indicating effective prompt collaboration. Adding the dynamic prompt yields gains of 13.5%, 13.26%, 8.4%/6.5% on all datasets, this is different from previous methods (Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021) where modalities don’t explicitly interact. Without the mapped prompt, fine-grained information from complementary modalities is lost, causing significant drops. The static prompt is also necessary, acting as a special expert always routed for capturing global-level features.
@@
Figure 4: More experts v.s. longer prompts. We compare increasing the number of experts, $k$, versus lengthen prompt, $l$. Expert-scaling consistently outperforms length-scaling, exhibiting a linear growth trend and do not deteriorate even when $k\to\infty$. Conversely, excessive prompt length has a negative impact on accuracy and speed.
@@
Expert-scaling is more expressive than length-scaling. To study the expressiveness of MoPE, we compare expert-scaling with length-scaling. Our starting point for MoPE is $l=6$ prompts and $k=2$ experts, which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer. We increase the number of $k$, and at each step we also report the result of increasing $l$ in vanilla prompt to the same total number of prompts. The results are summarized in Fig. 4.
@@
As illustrated in the figure, adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement which avoids over-length deterioration. In contrast, length-scaling does not lead to a linear performance improvement and is prone to performance deterioration, which is consistent with previous findings (Jia et al. 2022; Yang et al. 2022a; Khattak et al. 2023; Li and Liang 2021). Additionally, longer prompts in vanilla prompt tuning can exacerbate computational overhead due to the quadratic time complexity of self-attention. This is why we cannot experiment with $l>199$ for the other two methods. By contrast, MoPE scales by conditioning the dynamic prompt on more experts while maintaining a fixed prompt length during actual self-attention. As a result, our method maintains a nearly constant time complexity.
@@
Figure 5: Scaling performance with increased training data. We show the performance of our method and representative methods as we progressively increase the amount of training data, or “#shots”. Our method outperform other prompt fusion methods at all data scales.
@@
MoPE scales better with more training data. Previous prompting methods have been shown to not scale well with respect to increased training data (Gao et al. 2022; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021; Yang et al. 2022a). To assess the scalability of MoPE, we sub-sample the training set with different shots to simulate a range of low-shot to high-shot learning scenarios. We then train our method and other prompt-tuning and fine-tuning methods on the same subset of SNLI-VE and MM-IMDB.
@@
As Fig. 5 shows, our MoPE-based method demonstrates superior scalability compared to other prompt tuning methods. Specifically, we consistently match the results of the fine-tuning method, SeqFuse, on the SNLI-VE dataset, while surpassing it on the MM-IMDB dataset. By contrast, the methods based on vanilla prompts do not scale well with respect to increased training data, resulting in a consistent performance gap between our method and the compared methods. This performance disparity becomes more pronounced on larger datasets, such as when training with $10^{5}$ shots or using the full training set.
@@
Frozen routing embedding outperform learned embedding. We ablate the effect of using frozen routing embeddings. As Tab. 4 shows, fixed routing embedding could achieve the same result as the trained one.
Routing Embed SNLI-VE $\uparrow$ UPMC Food $\uparrow$ MM-IMDB $\uparrow$
@@
How MoPE could be more expressive. We have demonstrated that the proposed MoPE scales better with trainable parameters and data. This increased expressiveness comes from two factors. First, previous methods use a shared prompt for all instances and neglect multimodal interplay. By contrast, our dynamic prompt with MoPE is adaptive, and finding an instance-wise optimal prompt can lead to better results. Secondly, MoPE circumvents the challenge of optimizing a long prompt (Petrov, Torr, and Bibi 2023). By doing so, we shift the difficulty from learning a universal long prompt to learning a parameterized router with multiple short prompts. As a result, our empirical expressiveness is closer to the theoretical upper bound of prompt tuning.
Conclusion
In this paper, we propose an adaptive and expressive prompt-based method for fusing unimodal models. Our method involves decomposing the prompt for instance-adaptive prompt learning. To effectively scale up the prompt while avoiding performance deterioration, we introduce MoPE, which improves the expressiveness of prompt tuning. Extensive experiments demonstrate that our method is highly parameter-efficient and scales better with dataset size and prompt length. We believe MoPE could be an efficient fusion method for a wide array of multimodal tasks.
@@ Appendix A Additional Details on Experiments
Architecture Details
Prompt Vector. Our prompt implementation closely follows VPT (Jia et al. 2022). Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer (Jia et al. 2022). Dropout with $p=0.1$ is applied to all prompts. However, we do not use the reparameterization trick for prompts introduced in the original prompt tuning methods (Li and Liang 2021), as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs (i.e., $\mathbf{W}_{x}$, $\mathbf{W}_{y}$, $f_{m}(\cdot)$). For Transformer architectures that employ a window attention mechanism (e.g., Swin (Liu et al. 2021b)), we duplicate the same prompt and prepend it to all windows for self-attention calculation, following the approach in VPT (Jia et al. 2022).
Mapper. We learn a mapper to map representations from complementary modality $\mathbb{Y}$ to the embedding dimension of main modality $\mathbb{X}$. Generally speaking, the mapper is implemented as a two-layer MLP with a bottleneck design, which shares similarities with previous work (Li et al. 2023). In our experiments, we set the bottleneck dimension as half of the dimension of the complementary representation, i.e., $d_{bot}=\lceil d_{y}/2\rceil$. Then, we apply a batch normalization layer and a GeLU activation to obtain a bottleneck feature $\psi_{bot}\in\mathbb{R}_{d_{bot}}$.
After obtaining this bottleneck feature, we apply another linear layer to project it to the dimension of $d_{x}$. However, some Transformer architectures use inconsistent $d_{x}$ in different layers. For example, in Swin-b (Liu et al. 2021b), the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single linear transformation. To circumvent this, we instead learn four separate linear projections, each used to project the shared bottleneck feature to different embedding dimensions. As a result, there is one single down-sampling layer and multiple up-sampling layers. The design is illustrated in Fig 7.
@@
Dataset
UPMC Food-101 (Wang et al. 2015) serves as a comprehensive multimodal dataset designed for fine-grained recipe classification. The dataset contains 90,840 image-text pairs for 101 food classes. We follow previous methods (Li et al. 2023; Kiela et al. 2019) to create a validation set of 5000 samples.
MM-IMDB (Arevalo et al. 2017) is a multimodal movie classification dataset. It comprises 25,956 pairs of images and texts, each pair including a movie poster and a plot summary. The dataset supports multi-label genre classification across a spectrum of 23 genres with highly imbalanced classes.
SNLI-VE (Xie et al. 2019) is a large-scale multimodal dataset with 565,286 image-text pairs. The task for this dataset is visual entailment, which means that the model should decide whether a hypothesis matches the given premise or not. This dataset provide image and text premise, while the hypothesis is always in text modality.
MUsTARD (Castro et al. 2019) is a dataset for multimodal sarcasm detection. The original dataset contains 690 video clips in MP4 format with an even number of sarcastic and non-sarcastic labels. Each video is provided with annotations in the language modality, describing the speaker and sentence in that video. Collectively, there are three modalities available: video, audio extracted from the video, and the language modality.
RefCOCO and RefCOCO+ (Kazemzadeh et al. 2014) are two datasets for referring image segmentation. Both datasets are built based on MSCOCO and contain around 19,900 images with 50,000 objects and 140,000 referring expressions. RefCOCO allows referring expressions of any type, while RefCOCO+ features expressions that do not contain object positions.
@@
For the MuSTARD dataset, the overall processing pipeline follows PromptFuse (Liu et al. 2022). For the audio modality, we extract WAV audio from the video using a sampling rate of 16,000 Hz. We also remove background noise using the Librosa package. For a batch of input, we zero-pad the WAV input to the same length. As for the video, while PromptFuse (Liu et al. 2022) uses a face detection model to only sample frames where the speaker’s face is visible, we evenly sample $8$ frames for all videos in time. The sampled frames are resized to $(224,224)$ for encoding. For the text modality, we use the speaker-dependent setting, which means that the speaker’s name is visible to the model during both training and inference. To achieve this, we simply concatenate the speaker’s name to the input utterance string.

### v3

@@ 4.2 Experiment Setups
P-SeqFuse. Similar as SeqFuse but prompt-tune each encoders. This baseline is comparable to CoCoOp [42] but with additional static prompts.
In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT [15], Frozen [34], PromptFuse and BlindPrompt [21], and PMF [20]. Among these methods, the setting in PromptFuse [21] is the most similar to ours, as it also assumes one modality to be a black-box and is by design compatible with more than two modalities. To ensure a fair comparison, all prompt-tuning baselines utilize the same prompt length ($l=6$) as our method.
Prompt SNLI-VE Acc($\uparrow$) UPMC Food Acc($\uparrow$) MM-IMDB F1($\uparrow$)
@@
[…]
@@
4.3 Quantitative Comparisons
Multimodal Classification. The quantitative results of multimodal classification with all methods are summarized in Tab 1. We also list the inference time (in milliseconds) and number of trainable parameters (in million) for each method.
Our method outperforms all prompt-based fusion methods, and is competitive with fine-tuning. Specifically, when compared with the fine-tuning baselines, SeqFuse and LateConcat, our method delivers competitive accuracy on the UPMC Food-101 dataset and superior results on the SNLI-VE, MM-IMDB and MUStARD datasets, while requiring as few as 0.8% of the trainable parameters.
The proposed method also outperforms all existing prompt fusion methods, including PromptFuse [21], BlindPrompt [21], and PMF [20]. Notably, our method ($k=4$) outperforms the current SOTA, PMF, by a significant margin on all datasets, while requiring $37\%$ fewer parameters. Moreover, our method is built upon weak assumption than PMF, which assumes both encoders share the same Transformer architecture with an ad-hoc interaction layer. Their method is hard be extended for the task with three modalities, such as MUStARD. Extending PMF to heterogeneous model architectures is also are challenge. By contrast, our method is compatible with more than two modality and different architecture.
Referring Expression Segmentation. Previous prompt fusion methods mainly test their performance on classification [20, 21, 34]. The improved expressiveness of MoPE allows learning of dense prediction tasks, for example, referring image segmentation. The result is summarized in Tab. 2. Our proposed method achieves the best result among all compared methods by a significant margin. In particular the proposed method even outperform the fine-tuning baseline SeqFuse using only 15.3% parameters. This experiment further demonstrate the improved expressiveness of proposed method.
4.4 Qualitative visualization of MoPE Routing
Through training, prompt experts in MoPE spontaneously specialize, and we present two examples in Fig. 3. We visualize the expert with the highest score as the routed expert for each instance pair. In the provided examples, we observe that expert-4 specialize in children-related concepts, while expert-12 focuses on crowds. These examples also demonstrate the interpretability of MoPE, as opposed to black-box conditional prompt such as CoOp [43]. More visualizations could be found in supplementary material.
5 Analysis and Discussion
@@
Our decomposed prompts are collaborative. We ablate each type of prompt in our instance-wise adaptive prompt decomposition, with results across 3 random seeds reported in Tab. 3. Our full method achieves the best performance, indicating effective prompt collaboration. Adding the dynamic prompt yields gains of 13.5%, 13.26%, 8.4%/6.5% on all datasets, this is fundamentally different from previous methods [21, 34] where modalities don’t explicitly interact. Without the mapped prompt, fine-grained information from complementary modalities is lost, causing significant drops. The static prompt is also necessary, it could be interpreted as a special expert always routed to capture global-level features. Adding this prompt allows the other experts to focus more on capturing instance-level features, leading to a slight performance gain.
@@
Figure 4: More experts v.s. longer prompts. We compare increasing the number of experts, $k$, versus lengthen prompt, $l$. Expert-scaling consistently outperforms length-scaling, exhibiting a linear growth trend. Conversely, length-scaling suffer from deterioration.
@@
Expert-scaling is more expressive than length-scaling. We compare expert-scaling with length-scaling. Our starting point for MoPE is $l=6$ prompts and $k=2$ experts (when $k\leq 1$, the MoPE degenerates into vanilla prompt), which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer. We increase the number of $k$, and at each step we also report the result of increasing $l$ in vanilla prompt to the same total number of prompts. The results are summarized in Fig. 4.
@@
Adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement that avoids over-length deterioration. In contrast, length-scaling does not lead to a linear performance improvement and is prone to performance deterioration, which is consistent with previous findings [11, 40, 14, 19]. Additionally, longer prompts in vanilla prompt tuning can exacerbate computational overhead due to the quadratic time complexity of self-attention. By contrast, MoPE scales by conditioning the dynamic prompt on more experts. As a result, we scale up the empirical expressiveness while maintaining a fixed prompt length during actual self-attention. As a result, our method maintains a nearly constant time complexity.
@@
Figure 5: Scaling performance with increased training data. We show the performance of our method and representative methods as we progressively increase the amount of training data, or “#shots”. Our method outperform other prompt fusion methods at all data scales. The performance gap between the best compared prompt fusion method is shaded.
@@
MoPE scales better with more training data. Previous prompting methods have been shown to not scale well with respect to increased training data [9, 21, 34, 40]. To assess the scalability of MoPE, we sub-sample the training set with different shots to simulate a range of low-shot to high-shot learning scenarios. We then train our method and other prompt-tuning and fine-tuning methods on the same subset of SNLI-VE and MM-IMDB.
@@
Our MoPE-based method demonstrates superior scalability compared to other prompt tuning methods, as demonstrated in Fig. 5. Specifically, we consistently match the results of the fine-tuning method, SeqFuse, on the SNLI-VE dataset, while surpassing it on the MM-IMDB dataset. By contrast, the methods based on vanilla prompts, PromptFuse and P-SeqFuse are less scalable with respect to increased training data, resulting in a consistent performance gap between our method and the compared prompt fusion methods. This performance disparity becomes more pronounced on larger datasets, such as when training with $10^{5}$ shots or using the full training set. This result demonstrates the effect of proposed MoPE in scaling up the expressiveness of prompt fusion.
@@
Our prompt fusion method is highly modular. The proposed fusion method abstracts the complimentary modality as a representation, allowing high modality of both modalities. This is different from PMF [20], which is not modular. In particular, our method allow arbitrary models to be seamless plug-ins for multimodal fusion. Such modularity is at least threefold: model architecture, the pre-training scheme, and the specific transfer learning method of the model. We exemplify each in Tab. 4.
Architecture Pretraining Transfer MM-IMDB F1($\uparrow$) BoW Bert* Fine-tuning 48.20/57.50 Transformer Bert [5] Frozen 58.86/66.13 Transformer GPT-2 [30] Frozen 34.03/50.84 Transformer Bert [5] Fine-tuning 60.34/67.27
Table 4: Our prompt tuning method are highly modular. We offer flexibility for complementary modality in at least three dimensions: model architecture, pretraining scheme, as well as transfer learning technique. (*): Bag-of-word initialized with Bert word embeddings.
Frozen routing embedding outperform learned embedding. We ablate the effect of using frozen routing embeddings. As Tab. 5 shows, fixed routing embeddings could slightly better performance, while not requiring training.
Routing Embd SNLI-VE Acc($\uparrow$) UPMC Acc($\uparrow$) Food $\uparrow$ MM-IMDB F1($\uparrow$)
@@
How MoPE could be more expressive. We have demonstrated that the proposed MoPE scales better with trainable parameters and data. This increased expressiveness comes from two factors. First, previous methods use a shared prompt for all instances and neglect multimodal interplay. By contrast, our dynamic prompt with MoPE is adaptive, and finding an instance-wise optimal prompt can lead to better results. Secondly, MoPE circumvents the challenge of optimizing a long prompt [28]. By doing so, we shift the difficulty from learning a universal long prompt to learning a parameterized router with multiple short prompts. As a result, our empirical expressiveness is closer to the theoretical upper bound of prompt tuning.
Limitations and future works. Despite improving empirical expressiveness, our prompt still functions in the same way as the vanilla prompts during the self-attention. This means that our method is closer to the theoretical upper-bound of expressiveness for prompt tuning but not surpass it. This is also reflected by the small gap between ours and fine-tuning in large datasets such as SNLI-VE and UPMC Food-101. Future work could focus on augmenting the way that prompt attends to the token embedding to scale up its theoretical expressiveness.
6 Conclusion
In this paper, we propose an method to mitigate the issue of lacking adaptiveness and expressiveness in existing prompt-based multimodal fusion method. Our method involves decomposing the unified long prompt for instance-adaptive prompt learning. To effectively scale up the prompt while avoiding performance deterioration, we introduce MoPE, which improves the expressiveness of prompt tuning. Extensive experiments demonstrate that our method is highly parameter-efficient and scales better with dataset size and prompt length. We believe MoPE could be an efficient fusion method for a wide array of downstream multimodal tasks.
@@ Appendix A Additional Details on Experiments
A.1 Architecture Details
Prompt Vector. Our prompt implementation closely follows VPT [11]. Specifically, for static prompts and prompt experts, we use uniform initialization $\mathcal{U}\sim(-\eta,\eta)$, where $\eta$ is calculated according to the embedding dimension and patch size of the Transformer [11]. Dropout with $p=0.1$ is applied to all prompts. However, we do not use the reparameterization trick for prompts introduced in the original prompt tuning methods [19], as the gradients of our dynamic prompt and mapped prompt are already rectified by MLPs (i.e., $\mathbf{W}_{x}$, $\mathbf{W}_{y}$, $f_{m}(\cdot)$). For Transformer architectures that employ a window attention mechanism (e.g., Swin [25]), we duplicate the same prompt and prepend it to all windows for self-attention calculation, following the approach in VPT [11].
Mapper. We learn a mapper to map representations from complementary modality $\mathbb{Y}$ to the embedding dimension of main modality $\mathbb{X}$. Generally speaking, the mapper is implemented as a two-layer MLP with a bottleneck design, which shares similarities with previous work [20]. In our experiments, we set the bottleneck dimension as half of the dimension of the complementary representation, i.e., $d_{bot}=\lceil d_{y}/2\rceil$. Then, we apply a batch normalization layer and a GeLU activation to obtain a bottleneck feature $\psi_{bot}\in\mathbb{R}_{d_{bot}}$.
After obtaining this bottleneck feature, we apply another linear layer to project it to the dimension of $d_{x}$. However, some Transformer architectures use inconsistent $d_{x}$ in different layers. For example, in Swin-b [25], the embedding dimension changes from $[128,256,512,1024]$, making it challenging to fit a single linear transformation. To circumvent this, we instead learn four separate linear projections, each used to project the shared bottleneck feature to different embedding dimensions. As a result, there is one single down-sampling layer and multiple up-sampling layers. The design is illustrated in Fig 7.
@@ A.1 Architecture Details
A.2 Dataset
UPMC Food-101 [36] serves as a comprehensive multimodal dataset designed for fine-grained recipe classification. The dataset contains 90,840 image-text pairs for 101 food classes. We follow previous methods [20, 15] to create a validation set of 5000 samples.
MM-IMDB [1] is a multimodal movie classification dataset. It comprises 25,956 pairs of images and texts, each pair including a movie poster and a plot summary. The dataset supports multi-label genre classification across a spectrum of 23 genres with highly imbalanced classes.
SNLI-VE [39] is a large-scale multimodal dataset with 565,286 image-text pairs. The task for this dataset is visual entailment, which means that the model should decide whether a hypothesis matches the given premise or not. This dataset provide image and text premise, while the hypothesis is always in text modality.
MUsTARD [4] is a dataset for multimodal sarcasm detection. The original dataset contains 690 video clips in MP4 format with an even number of sarcastic and non-sarcastic labels. Each video is provided with annotations in the language modality, describing the speaker and sentence in that video. Collectively, there are three modalities available: video, audio extracted from the video, and the language modality.
RefCOCO and RefCOCO+ [13] are two datasets for referring image segmentation. Both datasets are built based on MSCOCO and contain around 19,900 images with 50,000 objects and 140,000 referring expressions. RefCOCO allows referring expressions of any type, while RefCOCO+ features expressions that do not contain object positions.
@@ A.3 Data Processing
For the MuSTARD dataset, the overall processing pipeline follows PromptFuse [23]. For the audio modality, we extract WAV audio from the video using a sampling rate of 16,000 Hz. We also remove background noise using the Librosa package. For a batch of input, we zero-pad the WAV input to the same length. As for the video, while PromptFuse [23] uses a face detection model to only sample frames where the speaker’s face is visible, we evenly sample $8$ frames for all videos in time. The sampled frames are resized to $(224,224)$ for encoding. For the text modality, we use the speaker-dependent setting, which means that the speaker’s name is visible to the model during both training and inference. To achieve this, we simply concatenate the speaker’s name to the input utterance string.